OpenStudy (anonymous):

What are the exact solutions of x2 - x - 4 = 0?

OpenStudy (anonymous):

do u have answer choices or what

OpenStudy (anonymous):

x2-x-4=0 divide by 2 on both sides x-x-4=0 +4 +4 ---------- x-x=4 it could be x^2=4 i believe

OpenStudy (anonymous):

\(x^2-x=4\) \(x^2-x+\frac{1}{4}=4+\frac{1}{4}\) \((x-\frac{1}{2})^2=\frac{17}{4}\) \( x-\frac{1}{2}=\pm\frac{\sqrt{17}}{2}\) \(x=\frac{1}{2}\pm \frac{\sqrt{17}}{2}\)

OpenStudy (anonymous):

Whenever you have everything factored equal to 0, then you take each part and make it equal 0. In this case: x - 2 = 0 x = 2 and x + 4 = 0 x = -4. The second way you could do it is to FOIL out the equation and put it into the quadratic formula (which would take longer, but if you want to try it, go ahead). Hope this helped (^_^).

OpenStudy (anonymous):

The exact solutions are x = 1/2 (1-sqrt(17)) ~~ -1.56155 and x = 1/2 (1+sqrt(17)) ~~ 2.56155.